# MARFT: Multi-Agent Reinforcement Fine-Tuning

### v2

@@ 8 Conclusion
In this article, we have presented Multi-Agent Reinforcement Fine-Tuning (MARFT), a novel framework that integrates RFT, LaMAS, and MARL to advance the capabilities of LaMAS systems. MARFT addresses several fundamental challenges in adapting conventional MARL to LaMAS, including asynchronous agent interactions, profile-aware architecture design, heterogeneity across agents, and the dynamic organization of multi-agent systems. Through the fusion of RFT with sequential decision-making paradigms, trust-region optimization techniques, and token-level adaptation mechanisms, MARFT enables efficient coordination and policy refinement while preserving the foundational capabilities of pretrained LLMs.
Preliminary experimental results on mathematical problem-solving tasks illustrate MARFT’s potential to surpass traditional single-agent strategies, demonstrating improved accuracy through emergent swarm intelligence, even in the presence of training instabilities. These findings suggest that MARFT holds significant promise for agentic tasks requiring complex, well-organized, distributed reasoning and acting.
The framework’s intrinsic scalability, privacy-preserving characteristics, and alignment with decentralized system architectures make it a compelling foundation for real-world deployment in domains ranging from logistics and collaborative robotics to blockchain-based ecosystems. Nevertheless, numerous theoretical and practical challenges remain. Key open problems include the development of dynamic environment designs with agentic tasks tailored for LaMAS, improvement of sample efficiency, and the absence of standardized and well-established frameworks for MARFT and a unified multi-agent communication protocol.
@@ Appendix A Detailed Framework Code Explanation
Abstract action-level PPO training code in the trainer:
@@
Then, for token-level training, we just need to modify the policy loss computation to take token logits as input.
@@ Appendix B Experiment Details
B.1 Hyperparameter Configurations for Preliminary Experiments
In Table 3, we provide the hyperparameter settings used in our preliminary experiments in Section 5. Recent experiments suggest that MARFT is highly sensitive to hyperparameter choices, and there may exist better configurations than those presented here.
@@ B.1 Hyperparameter Configurations for Preliminary Experiments
rollout threads 4 num mini-batch 1 ppo epoch 1
optimizer Adam hidden size 64 use huber loss True
optim eps 1e-5 critic loss coef 1 entropy coef 0.1
@@
context window 2048 max new tokens 512 clip param 0.2
Table 3: Hyperparameter configuration for the preliminary experiments.
@@
rollout threads 4 num mini-batch 1 ppo epoch 1
optimizer Adam hidden size 64 use huber loss True
optim eps 1e-5 critic loss coef 1 entropy coef 0
@@
context window 2048 max new tokens 512 clip param 0.2
Table 4: Hyperparameter configuration for experiment curve in Figure 4(a).
B.2 More Preliminary Experimental Results
We are concurrently conducting experiments on MATH(Hendrycks et al., 2021) and other environments. More correctness curves during the tuning process are demonstrated in Figure 4.
(a) Correctness of dual LaMAS with hyperparameter configuration shown in Table 4.
(b) Correctness of dual LaMAS with hyperparameter configuration shown in Table 5.
(c) Correctness of single agent with hyperparameter configuration shown in Table 5.
Figure 4: Training curves of more preliminary experiments.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
MAS learning rate 1e-6 critic learning rate 1e-5 use gae True
rollout threads 4 num mini-batch 1 ppo epoch 1
optimizer Adam hidden size 64 use huber loss True
optim eps 1e-5 critic loss coef 1 entropy coef 0.01
gamma 0.99 gae gamma 0.95 max grad norm 0.5
context window 2048 max new tokens 512 clip param 0.2
Table 5: Hyperparameter configuration for experiment curve in Figure 4(b) and 4(c).
@@ B.3 Profiles
Single-agent profile:

### v3

@@ 8 Conclusion
In this article, we propose a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT), together with a brand-new POMDP called Flex-POMDP to better align with the LaMAS optimization in realistic scenarios, and present a universal algorithmic framework, MARFT, that integrates RFT, LaMAS, and MARL to advance the capabilities of LaMAS systems. MARFT addresses several fundamental challenges in adapting conventional MARL to LaMAS, including asynchronous agent interactions, profile-aware architecture design, heterogeneity across agents, and the dynamic organization of multi-agent systems. Through the fusion of RFT with sequential decision-making paradigms, trust-region optimization techniques, and token-level adaptation mechanisms, MARFT enables efficient coordination and policy refinement while preserving the foundational capabilities of pretrained LLMs.
Preliminary experimental results on mathematical problem-solving tasks illustrate the potential of MARFT to surpass traditional single-agent strategies, demonstrating improved accuracy through emergent swarm intelligence, even in the presence of training instabilities. These findings suggest that MARFT holds significant promise for agentic tasks requiring complex, well-organized, distributed reasoning and acting.
The paradigm’s intrinsic scalability, privacy-preserving characteristics, and alignment with decentralized system architectures make it a compelling foundation for real-world deployment in domains ranging from logistics and collaborative robotics to blockchain-based ecosystems. Nevertheless, numerous theoretical and practical challenges remain. Key open problems include the development of dynamic environment designs with agentic tasks tailored for LaMAS, improvement of sample efficiency, and the absence of standardized and well-established frameworks for MARFT and a unified multi-agent communication protocol.
@@ Appendix A Detailed Framework Code Explanation
Abstract action-level MARFT training code in the trainer. (For token-level training, we just need to modify the policy loss computation to take token logits as input):
@@ Appendix B Experiment Details
B.1 More Experiment Curves
(a) Average Step Reward (MATH).
(b) Average Step Reward (CMATH).
Figure 7: Curves of average step reward during Duo MARFT with different optimization granularity (computed over 3 random seeds).
(a) Episodic Return (MATH).
(b) Episodic Return (CMATH).
Figure 8: Curves of episodic return during Solo MARFT (computed over 3 random seeds).
(a) Episodic Return (MATH).
(b) Episodic Return (CMATH).
(c) Episodic Return (MATH).
(d) Episodic Return (CMATH).
Figure 9: Curves of episodic return during Duo MARFT with and without warm-up. The warm-up steps is set to 1k.
B.2 Hyperparameter Configurations for Experiments
Table 5: Hyperparameter configuration for learning dynamics in Figure 4.
@@ B.2 Hyperparameter Configurations for Experiments
rollout threads 1 num mini-batch 1 ppo epoch 1
max horizon 2 hidden size 64 episode length 8
optim eps 1e-5 critic loss coef 1 entropy coef 1e-3
@@
context window 2456 max new tokens 512 clip param 0.2
Table 6: Hyperparameter configuration for token-level MARFT in Figure 5.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
MAS learning rate 3e-6 critic learning rate 1e-5 use gae True
rollout threads 1 num mini-batch 1 ppo epoch 1
max horizon 2 hidden size 64 episode length 8
optim eps 1e-5 critic loss coef 1 entropy coef 1e-3
gamma 0.99 gae gamma 0.95 max grad norm 0.5
context window 2456 max new tokens 512 clip param 0.2
Table 7: Hyperparameter configuration for tuning Solo in Table 4.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
MAS learning rate 3e-6 critic learning rate 1e-5 use gae True
rollout threads 2 num mini-batch 1 ppo epoch 1
max horizon 2 hidden size 64 episode length 8
optim eps 1e-5 critic loss coef 1 entropy coef 1e-3
gamma 0.99 gae gamma 0.95 max grad norm 0.5
context window 1536 max new tokens 512 clip param 0.2
Table 8: Hyperparameter configuration for warm-up ablation in Figure 6.
@@
rollout threads 1 num mini-batch 1 ppo epoch 1
max horizon 2 hidden size 64 episode length 8
optim eps 1e-5 critic loss coef 1 entropy coef 1e-3
@@
context window 2456 max new tokens 512 warmup steps 5000
@@ B.3 Profiles
Solo (Single-agent) profile: